OpenStudy (anonymous):

Give an example of a rational function that has a horizontal asymptote at y = 1 and a vertical asymptote at x = 4.

OpenStudy (zzr0ck3r):

\(\frac{ax+2}{a(x-1)}\)

OpenStudy (zzr0ck3r):

pick anything for a....

OpenStudy (imstuck):

\[\frac{ x+1 }{ x-4 }\]

OpenStudy (zzr0ck3r):

wow im out of it for some reason I read x=1...

OpenStudy (anonymous):

\[(1/(x-1))+4\] \[\frac{ 1 }{ x-1 } + 4\]

OpenStudy (anonymous):

so catch.me is correct?

OpenStudy (anonymous):

graph it at http://rechneronline.de/function-graphs/

OpenStudy (zzr0ck3r):

everyone is

OpenStudy (zzr0ck3r):

except me lol

OpenStudy (anonymous):

okay lol thanks everyone for your help

OpenStudy (anonymous):

LOL you read it wrong.It sometimes happen with me :D

OpenStudy (zzr0ck3r):

\(\huge :)\)